A group of 30 tourists paid a total of $686 in bus fares. If the adult fare was $28 and children paid half fare, how many adults were in the group?
Answer by pakhi(24)

Let the number of adults in the group be A.
Let the number of children in the group be C.
Therefore according to the problem
Condition 1 A child's fare was half of that of an adult's fare i.e. 28/2= $14.

Condition 2 Total number of tourists were equal to 30
A + C = 30 or C = 30 - A ------------------(1)
Condition 3 Total bus fare was $686 which was equal to the bus fares of A adults @ $28 each plus the fares of C children @ $14 each
28A + 14C = 686
Putting the value of C from equation 1 we get
28A + 14(30 - A) = 686
or 28A + 420 - 14A = 686
or 14A = 686 - 420 = 266
or A = 266/14 = 19
Also C= 30 - A = 30 - 19 - 11
So the number of adults in the bus was 19 and the number of children was 11.